500 g of water at 18 ° C and 0 g of 80 g of ice cubes are added. Dish own heat capacity is 100 J / K. Do you melt everything? If it melts, then what is the temperature of the final state, and if not melt, how much ice will remain without melting?
1
Expert's answer
2019-03-18T12:49:52-0400

Check up if the heat lost by water and dish will be enough to melt all the ice.

Heat required to melt all the ice is

"q_1=m \Delta H_{fus}=80g\times334\frac{J}{g}=26720J"


Heat lost by water when temparature drops from"18^\circ C to 0^\circ C" is

"q_2=cm \Delta T=4.186\frac{J}{g\times^\circ C}\times500g\times(0^\circ C -18^\circ C)= -37674 J"

 

Heat lost by dich is

"q_3=c\times \Delta T = -100\frac{J}{^\circ C}\times (0^0C-18^0C) = -1800 J"


Total lost of heat

"q_2 +q_3 = -37674J -1800J = -39474 J"


We can see that "q_1 <q_2 + q_3"  then all the ice will melt.


Determine the temperature of the final state:

"q_{gained} =q_{lost}"

"q_{gained} = q_1 +q_4"

"q_1=m \Delta H_{fus}=80g\times334\frac{J}{g}=26720J"

Heat gained by 80 g of melted ice when the temperature raises from "0^\circ C to T_{final}"  is

"q_4 = cm \Delta T"

"q_4 = 4.186f\frac{J}{g\times^\circ C}\times80g\times(T_{final} -0^\circ C)"


"q_{lost} = -(q_2+q_3) = -(4.186\frac{J}{g\times^\circ C}\times500g\times(T_{final} -18^\circ C) + 100\frac{J}{^\circ C} \times(T_{final} - 18^\circ C)"


"q_{lost} =q_{gained}"

"26720 + 4.186 \frac{J}{g\times^\circ C}\times80g\times(T_{final} -0^\circ C)= -(4.186 \frac{J}{g\times^\circ C}\times500g\times(T_{final} -18^\circ C) + 100\frac{J}{^\circ C}\times(T_{final} - 18^\circ C))"


"T_{final} = 5.05 ^\circ C"


Answer: "T_{final} = 5.05 ^\circ C"